# Inter-turbine Modelling of Wind-Farm Power using Multi-task Learning

### v1

@@ 6.1 BSBR Model Demonstration
To first demonstrate probabilistic predictions that can be made with the BSBR models (with no-pooling), Figure 9 shows predictions from models of turbines on both the west and east-side of the wind-farm in subplots (a) and (b) respectively. Test data are indicated by red crosses, model mean predictions are indicated by circles (which are coloured by the value of the wind direction), and model uncertainty is represented by the blue shaded area–indicating the 95th-percentile uncertainty bounds.
@@
Figure 9: BSBR NP model predictions for two separate turbines. Test data are shown by red crosses, mean predictions are shown as circles, coloured by the wind direction. The blue shaded areas show the 95th-percentile of the predictive variance.
@@
There are a number of interesting observations. Firstly, the predictive uncertainty appears to match the data, with reduced variance at the lower and upper ends of the curve, and increased variance in the transitional region between turbine minimum and maximum power. Approximately 5% of the test points also lie outside the 95th-percentile of the predictive variance, as should be expected for a correctly-trained model. The “rough” edges of this uncertainty are from data points that, while having a very similar freestream wind speed, may have significantly different wind directions, which also contribute to the predictive uncertainty. Secondly, while the majority of the influence on output power appears to be associated with the freestream wind speed (as expected), the models have allocated some weight to the wind-directional features, resulting in the variance in the mean predictions for a given freestream wind speed; this is mostly prevalent in the transitional region between minimum and maximum power, and matches intuition since turbine wakes affect this region the most. However, in this region it is clear that significant deviation between model mean predictions and the test data remains. This deviation is likely the result of a number of factors, including the oversimplification of the wake effect by assuming it can be approximated based on the wind direction. In reality, this fails to account for atmospheric conditions such as wind shear, turbulence intensity, thermal stratification and air-density variations, which are known to have an impact on wake propagation [35, 36]. Furthermore, in calculating the input features, the wind direction (yaw angle) is assumed to be the same for all turbines - this is not the case, as there is a distribution of localised wind-directions across the wind-farm (yaw angles), which will likely change the wake patterns. There are also additional sources of noise within the data, such as sensor calibration errors in the yaw and wind speed measurements which are used in calculating the population-level features, as well as possible changes in the underlying behaviour of the turbines, perhaps because of wear or damage during the period of the training data used. Additionally, curtailed, shut-down or even power-boosted turbines upstream of a given turbine will further generate noise in the wake patterns.
@@
Despite these problems, comparison of the west and east-side models yields some evidence that a level of turbine-dependent wind-directionality has been captured. For example, for the west-side turbine, it is not expected to be influenced by wakes when wind approaches from the western direction (equivalent to a value of 0 for feature two). Conversely, the east-side turbine is not expected to be influenced by wakes when wind blows from the east (equivalent to a value of 1 for feature three). In each case, the models appear to predict higher levels of power when these conditions are true, and less power when they are not, as one would expect to be the case in reality. While not shown here, these characteristics can also be seen for other turbine models around the perimeter of the wind-farm, to similar or lesser degrees.
@@
To explore this model behaviour further, the learned model parameter vectors which result in differences in predicted values may be inspected. In this case, each of the vectors $\boldsymbol{\eta}$ and $\boldsymbol{\zeta}$ consist of 18 parameters, with parameters 1-6 associated with Feature One, 7-12 Feature Two and 13-18 Feature Three. Figure 12 shows the value of selected parameters from the $\boldsymbol{\eta}$ vector plotted for each turbine within the wind-farm. The colour of the points denotes the values of the parameters as indicated by the colour bars.
@@
For Parameter Four (associated with Feature One), there does not appear to be an obvious spatial correlation. However, while their magnitude is considerably smaller than Parameter Four, Parameters Eight and 13 (associated with Features Two and Three), shown in subfigures 11(a) and 12(a), show clear spatial correlation across the wind-farm. For Parameter Eight, there is an approximate negative correlation from west to east, whilst there is an approximate south to north negative correlation for Parameter 13. While not shown here, similar observations can be made for the remaining parameters in the $\boldsymbol{\eta}$ and $\boldsymbol{\zeta}$ vectors. These spatial differences in parameters represent the wake effect captured by the NP models, and motivated the development of the metamodel, which aims to remove the requirement for data from each turbine location, by utilising the correlation in parameters. The use of a partially-pooled hierarchical model in this case also allows “statistical strength” to be shared between turbine-specific parameters, via the population-level parameters.
@@
(a) 8th parameter in the learned $\boldsymbol{\eta}$ vectors.
@@ 6.2 Metamodel
The metamodel learns linear models over the parameter vectors, for turbine $x$ and $y$ coordinates. The right-hand plots in sub-Figures 13(a) and 14(a) show the metamodel mean predicted value by colour, plotted spatially, for $\boldsymbol{\eta}$ Parameters Four and Eight respectively. Filled circles denote the locations of the turbines used for training the metamodel, whilst unfilled circles represent the locations of the remaining unobserved turbines used in testing. These predictions are compared directly with the NP BSBR models, shown in the left-hand plots. As with Figure 10(a), there is limited spatial correlation for Parameter Four seen for the NP model, which for the metamodel resulted in very little spatial difference in this parameter across all turbines. In contrast, for Parameter Eight it can be seen that the metamodel has learned the more significant correlation seen in the NP model parameters. These observations could suggest that the linear model assumption offers a level of regularisation, and is able to capture intuitively the general observable spatial correlation in parameters.
@@
(a) 8th parameter in the learned $\boldsymbol{\eta}$ vectors.
Figure 14: Mean predicted values for the fourth (subplot (a)) and eighth (subplot (b)) entries in the $\boldsymbol{\eta}$ parameter vector, plotted spatially for each turbine in wind-farm Ciabatta. In each subplot, the left-hand plots show predictions from the NP model, whilst the right-hand plots show metamodel predictions. Filled circles represent turbines used in training, whilst unfilled circles represent turbines used in testing only. The colours of the points represent the values of the parameters.
@@ 6.2.1 Model Scores
The metamodel is compared against the benchmark NP, PP and CP models, using the scoring metrics. Figure 15(a) shows the NMSE per turbine, for each model, with a lower score indicating an improved performance on the turbine’s associated test data. Figure 15(b) shows the JLL per turbine, for each model, with a higher score indicating an improved performance on the turbine’s associated test data. In both figures, the vertical dashed lines denote turbines used in training.
(a) NMSE scores
(b) JLL scores
@@
Figure 15: Comparison of model scores for each turbine in wind-farm Ciabatta. For the metamodel, CP and PP models, the vertical dashed lines denote turbines used in training. Coloured dashed horizontal lines represent the mean score across all turbines for each model.
For both metrics, it can be seen that the CP model performs significantly worse than the other models, which perform more similarly to each another; this suggests a benefit in accounting for task-level differences, which the CP model does not. There are some instances however where the NP models perform worse the CP model on JLL, in particular for Turbines 17 and 58. During MCMC sampling, on some occasions a few of the NP model estimations resulted in divergences, which suggests the sampler was having difficulties exploring the posterior geometry; this was not the case for the other models, which may suggest that their sharing of information was helpful in parameter estimation [32].
More broadly, it can be seen on average that the PP model performs similarly but marginally better than the NP model, however the metamodel shows a more distinct improvement on both metrics, showing an average improvement of approximately 2.3% in NMSE and 60 in JLL compared to the PP model. Moreover, the metamodel nearly always scores the highest on both metrics across all turbines, (including the previously-unobserved testing turbines); this provides validation that the assumptions used in the construction of the metamodel provide valuable information in the spatial prediction of wind-turbine power.
@@
A final observation is that there are relatively-high NMSE scores across all models for Turbines 40–50 (and relatively low JLL), followed by lower scores of Turbines 50–60; to explore this further, the scoring metrics for the metamodel are plotted spatially in Figure 16. Here, in general the scores appear to be worse towards the lower end of the wind-farm; this matches intuition, since these turbines also follow a less rigid spatial pattern than the turbines at the upper end, which is likely to add additional noise to the spatial wake patterns of these turbines.
@@
Figure 16: Maps of model scores across the wind farm. The left-hand plot shows the NMSE per turbine, whilst the right-hand plot shows the JLL per turbine.
@@ 6.2.2 Predicting Unobserved Turbines
The models are now compared in prediction, for an unobserved turbine i.e. a turbine that was not used in training for the metamodel, PP and CP models. Figures 17(a)17(d) compare the learned linear models $f_{1}$ and $f_{2}$ across all four models. A gridded input over all three features was used, taking averages of the predictive mean and variance over the wind-directional Features Two and Three. Model mean predictions are shown as solid lines, while the shaded areas represent the $95^{th}$-percentile uncertainties. For the CP model, the predictive variance is relatively high for all values; which is reflective of the model not accounting for task-specific differences, and capturing the variance of all tasks in the learned parameters. The NP model also has relatively high predictive variance, particularly at the ends of the feature space where training data become more scarce; on the other hand, the PP and metamodels benefit from the sharing of statistical strength between turbines, and have reduced predictive variance across the feature space as a result, despite this turbine being unobserved.
@@
Figure 17: Posterior predictive distributions for functions $f_{1}$ and $f_{2}$, for turbine 35, using a gridded input over all three features. Averages were taken over Features Two and Three. Solid lines represent the mean prediction, while shaded areas show the $95^{th}$-percentile uncertainty.
These observations carry into the full probabilistic power-predictions, as shown in Figure 18. Whilst a typical power-curve is “sigmoidal” in shape, in this case the shape of the model predictions differ slightly at higher freestream wind speeds, as a result of the stratified sampling carried out on the training and test data; this results in relative data sparsity at maximum power, particularly beyond a freestream wind speed of 0.8. For the NP model, this sparsity results in relatively high predictive variance in this region, with the model mean prediction also being partially “dragged down” towards zero - the default behaviour of B-spline models with no data. On the other hand, the metamodel and PP models benefit from the sharing of statistical strength between model parameters, resulting in reduced predictive variance; this demonstrates the advantage of the hierarchical modelling approach, although it does come with the cost of additional computational expense.
@@
Figure 18: Probabilistic power-predictions on the “unobserved” turbine 35, over the gridded input data.

### v2

@@ 6.1 BSBR Model Demonstration
To first demonstrate probabilistic predictions that can be made with the BSBR models (with no-pooling), Figure 10 shows predictions from models of turbines on both the west and east-side of the wind-farm in subplots (a) and (b) respectively, using the 1000 testing data points in each case. Test data are indicated by red crosses, model mean predictions are indicated by circles (which are coloured by the value of the wind-direction), and model uncertainty is represented by the blue shaded area–indicating the 95th-percentile of predictions.
@@
Figure 10: NP BSBR model predictions for two separate turbines. Test data (1000 points) are shown by red crosses, mean predictions are shown as circles, coloured by the wind-direction. The blue shaded areas show the 95th-percentile of the predictive variance.
@@
There are a number of interesting observations. Firstly, the predictive uncertainty appears to match the data, with reduced variance at the lower and upper ends of the curve, and increased variance in the transitional region between turbine minimum and maximum power. Approximately 5% of the test points also lie outside the 95th-percentile of the predictive variance, as should be expected for a correctly-trained model. The “rough” edges of this uncertainty are from data points that, while having a very similar freestream wind-speed, may have significantly different wind-directions, which also contribute to the predictive uncertainty. Secondly, while the majority of the influence on output power appears to be associated with the freestream wind-speed (as expected), the models have allocated some weight to the wind-directional features, resulting in the variance in the mean predictions for a given freestream wind-speed; this is mostly prevalent in the transitional region between minimum and maximum power, and matches intuition since wind-turbine wakes affect this region the most. However, in this region it is clear that significant deviation between model mean predictions and the test data remains. This deviation is likely to be the result of a number of factors, including the oversimplification of the wake-effect in assuming it can be approximated based on the wind-direction. In reality, this fails to account for atmospheric conditions such as wind shear, turbulence intensity, thermal stratification and air-density variations, which are known to have an impact on wake propagation [36, 37]. Furthermore, in calculating the input features, the wind-direction (yaw angle) is assumed to be the same for all turbines—this is not the case, as there is a distribution of localised wind-directions across the wind-farm (yaw angles), which will likely change the wake patterns. There are also additional sources of noise within the data, such as sensor calibration errors in the yaw and wind-speed measurements which are used in calculating the population-level features, as well as possible changes in the underlying behaviour of the turbines, perhaps because of wear or damage during the period of the training data used. Additionally, curtailed, shut-down or even power-boosted turbines upstream of a given turbine will further generate noise in the wake patterns.
@@
Despite these problems, comparison of the west and east-side models yields some evidence that a level of turbine-dependent wind-directionality has been captured. For example, for the west-side turbine, it is not expected to be influenced by wakes when wind approaches from the western direction. Conversely, the east-side turbine is not expected to be influenced by wakes when wind blows from the east. In each case, the models appear to predict higher levels of power when these conditions are true, and less power when they are not, as one would expect to be the case in reality. While not shown here, these characteristics can also be seen for other turbine models around the perimeter of the wind-farm, to similar or lesser degrees.
@@
To explore this model behaviour further, the learned model parameter vectors which result in differences in predicted power may be inspected. In this case, each of the vectors $\boldsymbol{\eta}$ and $\boldsymbol{\zeta}$ consist of 18 parameters, with parameters 1-6 associated with Feature One, 7-12 Feature Two and 13-18 Feature Three. Figure 13 shows the value of selected parameters from the $\boldsymbol{\eta}$ vector plotted for each turbine within the wind-farm. The colour of the points denotes the values of the parameters as indicated by the colour bars.
@@
For Parameter Four (associated with Feature One), there does not appear to be an obvious spatial correlation. However, while their magnitude is smaller than Parameter Four, Parameters Ten and 13 (associated with Features Two and Three), shown in subfigures 12(a) and 13(a), show clear spatial correlation across the wind-farm. For Parameter Ten, there is an approximate negative correlation from east to west, while there is an approximate south to north negative correlation for Parameter 13. While not shown here, similar observations can be made for some of the remaining parameters in the $\boldsymbol{\eta}$ and $\boldsymbol{\zeta}$ vectors. These spatial differences in parameters represent the wake-effect captured by the NP models, and motivated the development of the metamodel, which aims to remove the requirement of having data from each individual turbine, by utilising these spatial correlations in the individual model parameters. The use of a hierarchical model in this case also allows “statistical strength” to be shared between turbine-specific parameters, via the population-level parameters.
@@
(a) 10th parameter in the learned $\boldsymbol{\eta}$ vectors.
@@ 6.2 Metamodel
The metamodel learns linear models over the parameter vectors, for turbine $x$ and $y$ coordinates. The right-hand plots in sub-Figures 14(a) and 15(a) show metamodel mean predicted values by colour, for $\boldsymbol{\eta}$ Parameters Four and Ten respectively, spatially. The black filled circles denote the locations of the turbines used for training the metamodel, while the unfilled black circles represent the locations of the remaining (unobserved) turbines that were also used in testing. These predictions are compared directly with the NP BSBR models, shown in the left-hand plots. As with Figure 11(a), there is no clear observable spatial correlation for Parameter Four seen for the NP model; there are also some potential outliers—such as those shaded more yellow in the upper half of the wind-farm. For the same parameter in the metamodel, the regularising effect of the linear model assumption removes the outliers in prediction, while perhaps only a weak correlation in learned values across the farm is observed. For Parameter Eight it can be seen that the metamodel has approximated the more clear observable correlation seen in the NP model parameters. These observations provide confirmation that the metamodel is working as intended.
@@
(a) 10th parameter in the learned $\boldsymbol{\eta}$ vectors.
Figure 15: Mean predicted values for the Fourth (subplot (a)) and Tenth (subplot (b)) entries in the $\boldsymbol{\eta}$ parameter vector, plotted spatially for each turbine in wind-farm Ciabatta. In each subplot, the left-hand plots show the mean predictions from the NP model, while the right-hand plots show the mean metamodel predictions. Filled circles represent turbines used in training, while unfilled circles represent turbines used in testing only. The points are shaded by their values.
@@ 6.2.1 Model Scores
The metamodel is compared against the benchmark NP, PP and CP models, using the scoring metrics. In the case of the NMSE, a lower score indicates improved model performance, while for the JLL a higher score indicates an improved model fit. Firstly, Figure 16(a) shows the NMSE per turbine, for each model, for the observed turbines only (i.e. those used in training). Similarly, Figure 16(b) shows the JLL per turbine, for each model. For both metrics, it can be seen that on average, the metamodel and PP models score similarly to each other, and marginally better than both the CP and NP models, which themselves score similarly to each another.
(a) NMSE scores on the observed turbines
(b) JLL scores on the observed turbines
@@
Figure 16: Comparison of model scores for each of the observed turbines in wind-farm Ciabatta. Dashed horizontal lines indicate mean scores per model.
Secondly, Figures 17(a) and 17(b) show the NMSE and JLL per turbine, for each model, for the unobserved turbines; as a reminder, the exception to this is the NP models, which require data from every turbine, and so all turbines are individually observed in this case. For these otherwise unobserved turbines, it can be seen that on average the metamodel performs marginally better than all three of the other models for both the NMSE and JLL metrics, including the NP models despite not having observed data from these turbines in training.
(a) NMSE scores on the unobserved turbines
(b) JLL scores on the unobserved turbines
Figure 17: Comparison of model scores for each of the unobserved turbines in wind-farm Ciabatta. Dashed horizontal lines indicate mean scores per model.
Across all observed and unobserved turbines, the CP model scores particularly poorly on the JLL metric in five instances: the observed turbines 16 and 19, and the unobserved turbines eight, 32 and 38. Since the CP model learns a single set of shared population-level parameters which are used for prediction, it is not able to adjust for turbine-specific differences, which may be the case in these instances. For the unobserved turbines, the PP model is limited in the same way as the CP model, since the shared population-level parameters must be used for making predictions; this explains their similar average scores. However unlike the CP model, the PP model does not share the same sharp drops in the JLL metric (turbines eight, 32 and 38), suggesting that partial-pooling of information provides improved adaptation to the model. In the case of the observed turbines, the PP model is able to use the learned turbine-specific weights in prediction, and so can make adjustments which improve the model scores relative to the CP model, which are of a similar performance to the metamodel.
The metamodel combines strengths from all three approaches: information is pooled reducing the effective data-scarcity (like the CP and PP models), whilst turbine-specific adjustments may also be made (like the PP and NP models). Importantly however, the metamodel can additionally make adjustments for unobserved tasks, resulting in a distinct improvement in model performance for these tasks. Compared to the next best performing model for the unobserved turbines (the NP models), the metamodel scores were improved by 3.2% and 9.8% for NMSE and JLL metrics respectively.
@@
A final observation is that there are some relatively-high NMSE scores (and reduced JLL scores) across all models for the observed turbines 12–14, and the unobserved turbines 32–35, compared to the surrounding turbines; to explore this further, the scoring metrics for the metamodel are plotted spatially in Figure 18. Here, in general the scores appear to be worse towards the lower end of the wind-farm; since these turbines follow a less rigid spatial pattern than the turbines at the upper end, they are likely to be a source of additional noise in the spatial wake patterns of these turbines, which may make accurate prediction of output power using just the freestream wind speed and direction features used in this work more challenging.
@@
Figure 18: Maps of the metamodel scores across the wind-farm. The left-hand plot shows the NMSE per turbine, while the right-hand plot shows the JLL per turbine.
@@ 6.2.2 Predicting Unobserved Turbines
The models are now compared in prediction, for an unobserved turbine—for consistency, the same turbine as previously predicted for in Figure 10(b) was chosen. Figures 19(a)19(d) first compare the learned linear models $f_{1}$ and $f_{2}$ across all four models. A gridded input over all three features was used, with averages of the predictive mean and $95^{th}$ percentile taken over the wind-directional features (Features Two and Three). Mean predictions are shown as solid lines, while $95^{th}$-percentiles are shown as shaded areas.
@@
Figure 19: Posterior predictive distributions for functions $f_{1}$ and $f_{2}$, for the unobserved turbine on the east-side of the farm, using a gridded input over all three features. Averages were taken over Features Two and Three. Solid lines represent the mean prediction, while shaded areas show the $95^{th}$-percentile.
To aid interpretation, $f_{1}$ is proportional to the mean prediction as per Equation (4), while $f_{2}$ is proportional to the precision (see Equation (5)), which is interpreted as the inverse of the variance. While the CP, PP and metamodels are broadly similar in these plots, there are significant differences in the function shape and $95^{th}$ percentile that can be seen for the NP model. As a result of the stratified sampling carried out on the training and test data, on a turbine-by-turbine basis there is relative data-scarcity at high freestream wind speeds; for an individual NP model such as this one, this scarcity is likely to result in both increased uncertainty in the trained model weights in this region of the curve, and a tendency for these weights to also be more biased towards their prior distribution, which in this case were centred on zero. These factors likely both contribute towards $f_{1}$ being pulled downwards at higher wind speeds (and with increased uncertainty), and $f_{2}$ (proportional to precision) being relatively smaller (resulting in increased variance in the beta distribution). However, for both $f_{1}$ and $f_{2}$ the CP, PP and metamodels are not completely free of the issues of data-scarcity, as at high freestream wind speeds (beyond 0.8), the curves begin to decrease, which does not match the shape of the power-curve training data—power does not decrease at high wind speeds, nor does its variance increase after rated power is reached (assuming the cut-out wind speed is not reached). This model behaviour is perhaps a limitation of its current design, which could be improved by placing tighter, more positive priors for the higher freestream wind speed weights, which could be reasonably justified based on known behaviour of wind-turbines. More generally for the $f_{1}$ and $f_{2}$ functions, it can be said that the metamodel, PP and CP models benefit from pooling of information, resulting in narrower $95^{th}$ percentile predictions for both functions, despite not having observed data from this turbine.
These observations can directly be seen in the full probabilistic power-predictions—shown in Figure 20. Here, the mean prediction of the NP model can be seen to be lower than the other models, with increased uncertainty at higher freestream wind speeds as indicated by the $f_{1}$ and $f_{2}$ functions. The CP, PP and metamodels all show broadly similar predictions, which are well-matched to the shape of power-curve data. In this space, the effects of the reducing $f_{1}$ and $f_{2}$ functions at high freestream wind speeds appears to be negligible.
@@
Figure 20: Probabilistic power-predictions for the unobserved turbine, over gridded input data, averaged for the wind-direction features
Finally, to demonstrate how the metamodel is able to make location-specific adaptations in prediction for different wind-directions (which the CP and PP models cannot), Figure 21 shows power predictions for the same unobserved turbine, for easterly ($90^{\circ}$), southerly ($180^{\circ}$), westerly ($270^{\circ}$) and northerly ($360^{\circ}$) winds. As was shown for the model predictions in Figure 10(b), the model mean prediction is greater in the wake-effected region of the power-curve during easterly winds (as compared to westerly winds), which matches intuition.
Figure 21: Probabilistic power-predictions for the metamodel for the unobserved turbine 35, over a gridded input of freestream wind-speeds, for easterly ($90^{\circ}$), southerly ($180^{\circ}$), westerly ($270^{\circ}$) and northerly ($360^{\circ}$) wind-directions.
6.3 Model Computational Complexity
To compare the computational complexity of the models, Table 2 compares both the computational complexity of the likelihoods, and the number of learned parameters in each case.
Model $T$ $B$ NN
Likelihood
complexity
Number of
parameters
No-pooling 80 18 250 $\mathcal{O}(NT)=\mathcal{O}(20000)$ $2B\times T=36\times 80=2880$
Complete-pooling 20 18 250 $\mathcal{O}(NT)=\mathcal{O}(5000)$ $2B=36$
Partial-pooling 20 18 250 $\mathcal{O}(NT)=\mathcal{O}(5000)$ $2BT+2B+2=758$
Meta model 20 18 250 $\mathcal{O}(NT)=\mathcal{O}(5000)$ $8B=144$
Table 2: Comparison of the model likelihood complexity, and the number of parameters for each model. $T$: number of training turbines, $B$: width of the design matrix, NN: data points per training turbine.
While the number of iterations, and the model architecture influence overall computational complexity, the likelihood calculation is repeated every HMC iteration, and so should contribute significantly to the overall computational cost; in this case, since the CP, PP and metamodels are trained on 20 of the turbines, their likelihood complexity is one quarter of the overall cost of the 80 NP models. Additionally, the total number of parameters learned in each model varied significantly; the NP and PP models learn parameters for every turbine used in training (totalling 2880 and 758 respectively), while the CP and metamodels only learn parameters at the population-level (36 and 144 respectively). A higher number of parameters are also likely to increase the per-iteration cost, while also increasing the complexity of the posterior geometry, which affects how efficiently HMC can move [31]. In practice, model training varied in the order of hours for the CP, PP and metamodels, while it was in the order of a day for all 80 of the NP models, which roughly correlates with the increased total likelihood complexity, and number of parameters for these models.